# Antisymmetric tensor influence on charged black hole lensing phenomena and time delay

### v1

@@ IV.1 Charged black hole influenced by an antissymetric tensor
In this section, it is pertinent to emphasize that we exclusively consider the solution without a cosmological constant. This restriction is due to the applicability of the Tsukamoto method tsukamoto2017deflection , which is strictly valid for asymptotically flat spacetimes. Having outlined the methodology above, we proceed to apply it to our metric described in Eq. (1). Thus, we obtain
$b_{c}=\frac{\sqrt{-\frac{\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)^{4}}{(l-1)^{2}\left((l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}\right)}}}{2\sqrt{2}}.$ (6)
@@
$\tilde{a}=\frac{\sqrt{-\frac{\left(\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+3(l-1)^{2}M\right)^{2}}{3(l-1)M\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+9(l-1)^{3}M^{2}+8Q^{2}}}}{\sqrt{2}},$ (7)
@@
$\begin{split}&\tilde{b}=\frac{\sqrt{-\frac{\left(\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+3(l-1)^{2}M\right)^{2}}{3(l-1)M\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+9(l-1)^{3}M^{2}+8Q^{2}}}}{\sqrt{2}}\\ &\times\ln\left(\frac{6(l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+16Q^{2}}{(l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}}\right)+I_{R}(r_{m})-\pi.\end{split}$ (8)
In contrast with what happens in the schwarschild and Reissen–Nordstrom cases, notice that the contribution to the parameter $\tilde{a}$ is fundamentally modified by the Lorentz violation $l$, as one should expect. Also, $I_{R}(r_{m})$ reads
$\begin{split}&I_{R}(r_{m})=\int_{0}^{1}\mathrm{d}z\left\{\frac{\sqrt{2}\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)}{l-1}\right.\\ &\left.\times\left[\frac{1}{\sqrt{-\frac{z^{2}\left(3(l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+8Q^{2}\right)}{(l-1)^{2}}}}\right.\right.\\ &\left.\left.-\frac{1}{\sqrt{\frac{z^{2}\left((l-1)M(2z-3)\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{3}M^{2}(2z-3)-2Q^{2}(z-2)^{2}\right)}{(l-1)^{2}}}}\right]\right\}\\ &=0.298565.\end{split}$ (9)
Here, we have computed $I_{R}(r_{m})$ numerically, with parameters set as $Q=M=1$ and $l=0.01$. Consequently, the deflection angle is expressed as
$\begin{split}a(b)=&+0.298565-\pi-\frac{\sqrt{-\frac{\left(\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+3(l-1)^{2}M\right)^{2}}{3(l-1)M\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+9(l-1)^{3}M^{2}+8Q^{2}}}}{\sqrt{2}}\\ &\times\ln\left[\frac{b}{\frac{\sqrt{-\frac{\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)^{4}}{(l-1)^{2}\left((l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}\right)}}}{2\sqrt{2}}}-1\right]\\ &+\frac{\sqrt{-\frac{\left(\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+3(l-1)^{2}M\right)^{2}}{3(l-1)M\sqrt{(l-1)\left(9(l-1)^{3}M^{2}+8Q^{2}\right)}+9(l-1)^{3}M^{2}+8Q^{2}}}}{\sqrt{2}}\\ &\times\ln\left(\frac{6(l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+16Q^{2}}{(l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}}\right)\\ &+\mathcal{O}\left\{\left(b-\frac{\sqrt{-\frac{\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)^{4}}{(l-1)^{2}\left((l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}\right)}}}{2\sqrt{2}}\right)\right.\\ &\left.\times\ln\left[b-\frac{\sqrt{-\frac{\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)^{4}}{(l-1)^{2}\left((l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}\right)}}}{2\sqrt{2}}\right]\right\}.\end{split}$ (10)
To facilitate understanding, we present in Fig. 2 the variation of the deflection angle with $b$ under various system conditions.
@@ V.1 Galactic phenomena: gravitational lensing by Sagittarius $A^{*}$
To analyze the observables, we adopt a distance of $D_{OL}=8.5$ Kpc genzel2010galactic and $l\sim 3.90708\times 10^{-12}$ based on literature junior2024spontaneous . With $b_{c}=\frac{1}{2\sqrt{2}}\sqrt{-\frac{\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)^{4}}{(l-1)^{2}\left((l-1)M\left(\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M\right)+2Q^{2}\right)}}$, we find $\theta_{\infty}\approx 25.64$ $\mu$arcsecs $+\mathcal{O}(l)$, where $\mathcal{O}(l)$ denotes first–order effects in the Lorentz violation parameter, approximately $10^{-12}$. For a better interpretation, Figs. 4 and 5 illustrate observables $s$ and $\tilde{r}$ with respect to the LV coefficient. Both figures employ logarithmic scales due to the small contribution of LV. Fig. 4 reveals that $s$ increases as $l$ rises, indicating greater separation between the first and subsequent relativistic images. Conversely, Fig. 5 demonstrates that the flux ratio of the first image decreases with increasing $l$.
Figure 4: Observable $\ln s$ for distinct values of $\ln l$.
Figure 5: Observable $\ln\tilde{r}$ for a variety of $\ln l$.
@@ VI Time delay analysis: theoretical framework
where $T_{0}=\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}+\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$ represents the time taken for light to propagate in the absence of a gravitational field. It is evident that the time delay $\Delta T$ increases continuously with the positions $r_{\text{S}}$ and $r_{\text{O}}$ of the light source and observer. To enhance clarity for the reader, we present Fig. 6, illustrating the time delay $\Delta T$ as a function of $r_{\text{S}}$ for the specified parameters $l=0.01$, $M=Q=r_{0}=b=1$, and $r_{\text{O}}=100$. A comparison with the Reissner–Nordström case is provided. Furthermore, $\Delta T$ is also shown as a function of $l$ and $Q$.
Figure 6: The time delay $\Delta T$ as a function of $r_{\text{S}}$ is evaluated for specific parameter values: $l=0.01$, $M=Q=r_{0}=b=1$, and $r_{\text{O}}=100$. A comparison with Reissner–Nordström is shown as well. $\Delta T$ is also shown as a function of $l$ and $Q$.
@@ VII Summary and conclusion
In this study, we investigated the gravitational lensing effects of a spherically symmetric charged black hole within the context of Lorentz violation, incorporating an antisymmetric tensor field known as the Kalb–Ramond field to address this symmetry-breaking effect. Our calculations were approached through two distinct methods: the weak and strong deflection limits. For the weak deflection limit, we utilized the Gauss–Bonnet theorem to derive our results. Conversely, in the strong deflection limit, we applied the Tsukamoto methodology, which provided us with measurable outcomes such as relativistic image positions and magnifications. Using these methodologies, we analyzed the gravitational lensing by Sagittarius $A^{*}$ and determined the corresponding observables as functions of the Lorentz violation parameter $l$. Finally, the time delay was also computed.
@@ Acknowledgments
A. A. Araújo Filho is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ), project No. 150891/2023-7. The author also wishes to express gratitude to Professor A. Ovgün for the insightful correspondence, and to N. Heidari for highlighting a specific reference that was crucial for the time delay calculation.

### v2

@@ IV.1 Charged black hole influenced by an antissymetric tensor
Now, particularly, let us focus on the metric (1) duan2023electrically :
$\mathrm{d}s^{2}=-\left(\frac{1}{1-l}-\frac{2M}{r}+\frac{Q^{2}}{(1-l)^{2}r^{2}}\right)\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{\left(\frac{1}{1-l}-\frac{2M}{r}+\frac{Q^{2}}{(1-l)^{2}r^{2}}\right)}+r^{2}\mathrm{d}\theta^{2}+r^{2}\sin^{2}\theta\mathrm{d}\varphi^{2}.$
In the asymptotic limit, the metric functions exhibit the following behavior: $\lim\limits_{r\to\infty}A(r)=\frac{1}{1-l}$, $\lim\limits_{r\to\infty}B(r)=1-l$, and $\lim\limits_{r\to\infty}C(r)=r^{2}$. At first glance, this appears to contradict the condition of asymptotic flatness imposed in the previous section. However, since $l$ is a constant, one can rescale the time and radial coordinates as $t\to t^{\prime}=\frac{t}{\sqrt{1-l}}$ and $r\to r^{\prime}=\sqrt{1-l}\,r$. Given that $l$ is small, representing the Lorentz violation coefficient, the metric in this limit simplifies accordingly
$\lim\limits_{r^{\prime}\to\infty}\mathrm{d}s^{2}=-\mathrm{d}t^{\prime 2}+\mathrm{d}r^{\prime 2}+\left(1-\frac{l}{2}+\mathcal{O}(l^{2})\right)r^{\prime 2}\left(\mathrm{d}\theta^{2}+\sin^{2}{\theta}\mathrm{d}\phi^{2}\right).$ (40)
Notably, this line element bears resemblance to that of a global monopole with a deficit solid angle $\delta=\frac{l}{2}$, accurate to first order in $l$ PhysRevLett.63.341 . As discussed in the previous section, due to the spherical symmetry, photon trajectories can be confined to the equatorial plane ($\theta=\pi/2$). Under this condition, the line element simplifies to:
$\lim\limits_{r^{\prime}\to\infty}\mathrm{d}s^{2}|_{\theta=\pi/2}=-\mathrm{d}t^{\prime 2}+\mathrm{d}r^{\prime 2}+\left(1-\frac{l}{2}+\mathcal{O}(l^{2})\right)r^{\prime 2}\mathrm{d}\phi^{2}.$ (41)
This spacetime can be interpreted as a cosmic string geometry intersected by planes of constant $z$. Consequently, the line element (41) is locally flat with a conical topology, and it is inherently not asymptotically flat due to the presence of a deficit angle $\delta=\frac{l}{2}$. At first glance, this might suggest that the general methodology outlined in the previous section is inapplicable. However, the requirement of asymptotically flat spacetimes is introduced primarily to ensure the positivity of the effective potential at infinity ($r\to\infty$), which guarantees the existence of photons at large distances (as discussed earlier and in tsukamoto2017deflection ). In this context, such a condition can be relaxed since $\lim\limits_{r\to\infty}V(r)=E^{2}>0$ is satisfied.
In this section, it is pertinent to emphasize that we exclusively consider the solution without a cosmological constant. This restriction is due to the applicability of the Tsukamoto method tsukamoto2017deflection . To the black hole under consideration, the corresponding photon sphere is written as heidari2024impact
$r_{ph}=\frac{\sqrt{9(l-1)^{4}M^{2}+8(l-1)Q^{2}}+3(l-1)^{2}M}{2(1-l)},$ (42)
and, by considering $l$ small, we obtain
$r^{(l\ll 0)}_{ph}\approx\frac{1}{2}\left(\sqrt{9M^{2}-8Q^{2}}+3M\right)+\frac{\left(-3M\sqrt{9M^{2}-8Q^{2}}-9M^{2}-4Q^{2}\right)l}{2\sqrt{9M^{2}-8Q^{2}}},$ (43)
where it gives rise to a natural restriction of the parameter $Q$ and $M$ which is
$9M^{2}-8Q^{2}>0,\quad\text{and}\quad 9M^{2}-8Q^{2}\neq 0,$ (44)
ensuring that the expression yields real and positive values. It is worth noting that the first term in Eq. (43) coincides with the result for the Reissner–Nordström black hole. In contrast, the second term arises purely as a correction introduced by the Lorentz–violating parameter $l$.
Having outlined the methodology above, we proceed to apply it to our metric described in Eq. (1). Thus, we obtain
$b_{c}=3\sqrt{3-3l}(1-l)M-\frac{\sqrt{3-3l}\,Q^{2}}{2\left((l-1)^{2}M\right)}.$ (45)
@@
$\tilde{a}=1+\frac{Q^{2}}{9M^{2}}+\left(\frac{5Q^{2}}{18M^{2}}-\frac{1}{2}\right)l.$ (46)
@@
$\begin{split}&\tilde{b}=\left[1+\frac{Q^{2}}{9M^{2}}+\left(\frac{5Q^{2}}{18M^{2}}-\frac{1}{2}\right)l\right]\left(\ln(6)-\frac{(3l+1)Q^{2}}{9M^{2}}\right)+I_{R}(r_{ph})-\pi.\end{split}$ (47)
In contrast to the Schwarzschild and Reissner–Nordström cases, the contribution to the parameter $\tilde{a}$ is significantly altered by the presence of Lorentz violation $l$, as expected. Also, $I_{R}(r_{ph})$ reads
$\begin{split}I_{R}(r_{ph})=\int_{0}^{1}\mathrm{d}z&\left(-\frac{1}{18M^{2}(3-2z)^{3/2}z}\right)\times\left\{-18(l-2)M^{2}(2z-3)\left(\sqrt{3}-\sqrt{3-2z}\right)\right.\\ &\left.-(5l+2)Q^{2}\left(3\sqrt{3}z^{2}-8\sqrt{3}z+4\sqrt{3-2z}z-6\sqrt{3-2z}+6\sqrt{3}\right)\right\}\\ &=\frac{(5l+2)Q^{2}\left(2\sqrt{3}-5+\ln(36)+\ln\left(7-4\sqrt{3}\right)\right)}{18M^{2}}-2(l-2)\ln\left(3-\sqrt{3}\right).\end{split}$ (48)
Consequently, the deflection angle present in Eq. (38) is expressed as
$\begin{split}a(b)=&-\left[1+\frac{Q^{2}}{9M^{2}}+\left(\frac{5Q^{2}}{18M^{2}}-\frac{1}{2}\right)l\right]\left[\frac{b}{3\sqrt{3-3l}\,(1-l)M-\frac{\sqrt{3-3l}Q^{2}}{2\left((l-1)^{2}M\right)}}-1\right]\\ &+\left[1+\frac{Q^{2}}{9M^{2}}+\left(\frac{5Q^{2}}{18M^{2}}-\frac{1}{2}\right)l\right]\left(\ln(6)-\frac{(3l+1)Q^{2}}{9M^{2}}\right)\\ &+\frac{(5l+2)Q^{2}\left(2\sqrt{3}-5+\ln(36)+\ln\left(7-4\sqrt{3}\right)\right)}{18M^{2}}-2(l-2)\ln\left(3-\sqrt{3}\right)-\pi\\ &\mathcal{O}\left[\left(b-3\sqrt{3-3l}(1-l)M-\frac{\sqrt{3-3l}\,Q^{2}}{2\left((l-1)^{2}M\right)}\right)\ln\left(b-3\sqrt{3-3l}(1-l)M-\frac{\sqrt{3-3l}\,Q^{2}}{2\left((l-1)^{2}M\right)}\right)\right].\end{split}$ (49)
To facilitate understanding, we present in Fig. 2 the variation of the deflection angle with $b$ under various system conditions. In the top–left panel, although the variations are subtle, an increase in the parameter $l$ results in a slight decrease in the magnitude of the deflection angle $a(b)$, considering $Q=0.1$ and $M=1$. In the top–right panel, keeping $Q=0.1$ and $l=0.01$ fixed, a larger mass $M$ leads to an increase in the magnitude of $a(b)$. Finally, the bottom panel reveals that raising the charge $Q$, while maintaining $M=1$ and $l=0.01$, causes a reduction in the deflection angle.
@@ V.1 Galactic phenomena: gravitational lensing by Sagittarius $A^{*}$
To analyze the observables, we adopt a distance of $D_{OL}=8.5$ Kpc genzel2010galactic and $l\sim 3.90708\times 10^{-12}$ based on literature junior2024spontaneous . With $b_{c}=3\sqrt{3-3l}(1-l)M-\frac{\sqrt{3-3l}\,Q^{2}}{2\left((l-1)^{2}M\right)}$, we find $\theta_{\infty}\approx 25.64$ $\mu$arcsecs $+\mathcal{O}(l)$, where $\mathcal{O}(l)$ denotes first–order effects in the Lorentz violation parameter, approximately $10^{-12}$. For a better interpretation, Figs. 4 and 5 illustrate observables $s$ and $\tilde{r}$ with respect to the LV coefficient. Both figures employ logarithmic scales due to the small contribution of LV. Fig. 4 reveals that $s$ increases as $l$ rises, indicating greater separation between the first and subsequent relativistic images. Conversely, Fig. 5 demonstrates that the flux ratio of the first image decreases with increasing $l$.
Figure 4: Observable $\ln s$ for distinct values of $\ln l$ and for $Q=0.1$.
Figure 5: Observable $\ln\tilde{r}$ for a variety of $\ln l$ and for $Q=0.1$.
@@ VI Time delay analysis: theoretical framework
where $T_{0}=\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}+\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$ represents the time taken for light to propagate in the absence of a gravitational field. If we assume that the parameters $Q$, $l$, and the impact parameter are small, the time delay can be expressed analytically as follows
$\begin{split}\Delta T=&\,2(l-1)r_{0}-\frac{8lM^{2}}{r_{0}-2M}+\frac{4lMQ^{2}}{(r_{0}-2M)^{2}}-\frac{2Q^{2}}{r_{0}-2M}+(1-l)r_{\text{O}}-\frac{2lMQ^{2}}{(r_{\text{O}}-2M)^{2}}\\ &+\frac{4lM^{2}}{r_{\text{O}}-2M}+\frac{Q^{2}}{r_{\text{O}}-2M}+(1-l)r_{\text{S}}+\frac{4lM^{2}}{r_{\text{S}}-2M}-\frac{2lMQ^{2}}{(r_{\text{S}}-2M)^{2}}+\frac{Q^{2}}{r_{\text{S}}-2M}\\ &+b^{2}\left(\frac{1}{r_{0}}-\frac{r_{\text{O}}+r_{\text{S}}}{2r_{\text{O}}r_{\text{S}}}\right)+2(2l-1)M\Big{[}2\ln(r_{0}-2M)-\ln(r_{\text{O}}-2M)-\ln(r_{\text{S}}-2M)\Big{]}\\ &-\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}-\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}.\end{split}$ (55)
Figs. 6 and 7 illustrate the behavior of the time delay $\Delta T$ as a function of the impact parameter $b$ and the parameters $Q$ and $l$, respectively. In Fig. 6, the left panel displays $\Delta T$ versus $b$ for various values of $Q$, indicating that an increase in the charge parameter $Q$ leads to a noticeable decrease in the magnitude of the time delay. Conversely, the right panel reveals that increasing the Lorentz–violating parameter $l$ also results in a reduction of $\Delta T$.
In Fig. 7, we further analyze the combined influence of $Q$ and $l$ on $\Delta T$. The left panel examines the variation of $\Delta T$ with $Q$ for different values of $l$, while the right panel explores the dependence on $l$ for fixed values of $Q$. Overall, the results suggest that increasing $Q$ tends to suppress the time delay across a broad range of $l$ values, whereas variations in $l$ yield only a slight reduction in $\Delta T$ for different values of $Q$.
Figure 6: The time delay $\Delta T$ is plotted as a function of the impact parameter $b$, with the left panel illustrating the behavior for various values of the charge $Q$, while the right panel displays the effect of the Lorentz–violating parameter $l$.
Figure 7: The time delay $\Delta T$ is presented as a function of the charge $Q$ for different values of the Lorentz–violating parameter $l$ in the left panel, while the right panel displays $\Delta T$ as a function of $l$ for several fixed values of $Q$.
@@ VII Summary and conclusion
This study investigated the gravitational lensing phenomena associated with a spherically symmetric charged black hole in the presence of Lorentz symmetry violation, modeled through an antisymmetric Kalb–Ramond tensor field.
The analysis was conducted within both the weak and strong deflection regimes. In the weak deflection limit, we employed the Gauss–Bonnet theorem to derive analytical expressions for the deflection angle $\alpha(b,l,Q)$. The results showed that, for increasing values of the Lorentz–violating parameter $l$ and the electric charge $Q$, the deflection angle decreased as a function of the impact parameter $b$. However, when the black hole mass $M$ was increased while keeping $l$ and $Q$ fixed, the magnitude of $\alpha(b,l,Q)$ increased.
In the strong deflection limit, we adopted the Tsukamoto method, which allowed us to compute key observables such as the positions and magnifications of relativistic images. Notably, these results were also derived analytically. We applied this framework to the supermassive black hole Sagittarius $A^{*}$ and evaluated the corresponding observables as functions of the Lorentz–violating parameter $l$. Within this approach, as $a(b)$ varied, both the charge $Q$ and the parameter $l$ contributed to a reduction in the deflection angle when the mass $M$ was fixed. In contrast, for fixed values of $Q$ and $l$, an increase in $M$ led to an enhancement of $a(b)$.
Lastly, we also computed the time delay. In general, as $b$ increased, the magnitude of $\Delta T$ decreased with increasing $Q$, for fixed $M$ and $l$. A similar trend was observed for increasing $l$ at fixed values of $M$ and $Q$.
@@ Acknowledgments
A. A. Araújo Filho is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ), project No. 150891/2023-7. Special thanks are extended to Ali Övgün and P. J. Porfírio for the insightful correspondence throughout the preparation and revision of the manuscript. The author also wishes to thank N. Heidari for highlighting a specific reference that was fundamental in the time delay calculation.
VIII Data Availability Statement
Data Availability Statement: No Data associated in the manuscript